# DNF-Avatar: Distilling Neural Fields for Real-time Animatable Avatar Relighting

### v1

@@ 1 Introduction
In recent years, researchers have proposed methods to create animatable human avatars using neural fields [50, 43] along with human body prior models [42]. The robustness of neural fields allows for the estimation of geometry and appearance properties from monocular videos. However, one drawback of these methods is the slow rendering speed due to the underlying neural radiance fields (NeRF [44]) representation and the use of physically based rendering (PBR). To achieve PBR, existing methods employ Monte Carlo ray tracing, which is accurate but usually requires tracing a large number of secondary rays to attain high-quality PBR results, whereas a typical NeRF model only requires tracing a single primary ray to render a pixel. Thus, even with various acceleration techniques for NeRF, the rendering of state-of-the-art relightable human avatars is still inefficient, taking several seconds to render a single frame [67, 70, 38, 9].
With the advent of 3D Gaussian splatting (3DGS [30] and follow-up 2DGS [21]), a bunch of works has shown that Gaussian splatting can achieve real-time rendering of human avatars when combined with human body prior models [53, 45, 49, 36, 31, 32, 20, 25, 81, 73, 40]. However, the majority of these works focus on the novel-view synthesis task and do not consider the relightable appearance.
@@ 2.3 Relightable Avatar
Typical approaches for human avatar relighting often reconstruct the intrinsic properties of humans via a multi-view capture system with controlled lighting [13, 62, 18, 3, 77, 24, 10, 56]. In the absence of multi-view data and known illumination, R4D [9] jointly recovers the geometry, material properties, and lighting using a NeRF-based representation [78, 51]. However, R4D conditions the NeRF representation on observation space encoding, making it hard to generalize to novel poses. RANA [23] train a mesh representation for multiple subjects with ground truth physical properties. Sun et al. [60] computes the shading color via spherical Gaussian approximations. However, those two methods do not model the visibility, leading to the lack of shadowing effects.
@@ 3.1 Teacher Model
[…]
@@ 3.2 Student Model
where $\boldsymbol{\mathbf{r}}$ is the camera ray, $\boldsymbol{\mathbf{u}}(\cdot)$ returns the $uv$ coordinate of the intersection point between the camera ray and the Gaussian primitives. $\hat{\mathcal{G}}(\cdot)$ is a bounded 2D Gaussian density function.
@@
[…]
@@
where $deg$ is the degree of SH. To model the shadowing effect caused by body articulation, we transform the point $\boldsymbol{\mathbf{\mu}}_{o}$ and the normal $\boldsymbol{\mathbf{n}}_{o}$ from the Gaussian primitive in observation space to the canonical space of each body part. Then, we query the $\hat{\text{AO}}^{p}$ for each part and multiply them together to get the final ambient occlusion:
@@ 4.4 Intrinsic Properties Comparisons
We also compare with R4D and IntrinsicAvatar for the task of intrinsic property decomposition on the RANA dataset. As shown in the Tab. 2, our method outperforms our teacher model in terms of normal consistency. This disparity becomes more visible in qualitative comparisons shown in Fig. 3: IA tends to have texture details baked into geometry, whereas our model successfully keeps meaningful wrinkles while discarding high-frequency noise during distillation, thanks to the smoothness prior from 2DGS. For albedo estimation, our student model achieves comparable accuracy to that of the teacher model. Although increasing the albedo distillation loss enables the student model to closely match the teacher’s albedo, this exact replication proves to be suboptimal for relighting under our approximated PBR pipeline for the student. Instead, our objective prioritizes improved relighting performance over precise intrinsic decomposition.
@@ 5 Conclusion
In this supplementary document, we provide additional materials to supplement our main submission. In the supplementary video, we show more relighting results using our method. The code will be made publicly available for research purposes.

### v2

@@ 1 Introduction
In recent years, researchers have proposed methods to create animatable [17, 29, 18, 77] and relightable [71, 76, 42, 9] human avatars using neural fields [54, 47] along with human body prior models [46]. The robustness of neural fields allows for the estimation of geometry and appearance properties from monocular videos. However, one drawback of these methods is the slow rendering speed due to the underlying neural radiance fields (NeRF [48]) representation and the use of physically based rendering (PBR). To achieve PBR, existing methods employ Monte Carlo ray tracing, which is accurate but usually requires tracing a large number of secondary rays to attain high-quality PBR results, whereas a typical NeRF model only requires tracing a single primary ray to render a pixel. Thus, even with various acceleration techniques for NeRF, the rendering of state-of-the-art relightable human avatars is still inefficient, taking several seconds to render a single frame [71, 76, 42, 9].
With the advent of 3D Gaussian splatting (3DGS [33] and follow-up 2DGS [23]), a bunch of works has shown that Gaussian splatting can achieve real-time rendering of human avatars when combined with human body prior models [57, 49, 53, 40, 34, 36, 22, 27, 90, 80, 44, 19, 89, 87]. However, the majority of these works focus on the novel-view synthesis task and do not consider the relightable appearance.
@@ 2.3 Relightable Avatar
Typical approaches for human avatar relighting often reconstruct the intrinsic properties of humans via a multi-view capture system with controlled lighting [13, 66, 20, 3, 84, 26, 10, 60, 72]. In the absence of multi-view data and known illumination, R4D [9] jointly recovers the geometry, material properties, and lighting using a NeRF-based representation [85, 55]. However, R4D conditions the NeRF representation on observation space encoding, making it hard to generalize to novel poses. RANA [25] train a mesh representation for multiple subjects with ground truth physical properties. Sun et al. [64] computes the shading color via spherical Gaussian approximations. However, those two methods do not model the visibility, leading to the lack of shadowing effects.
@@ 3.1 Teacher Model
$T(\mathbf{x},\mathbf{y})$ $=\text{exp}\left(-\int_{0}^{\|\mathbf{y}-\mathbf{x}\|}\sigma_{t}(\mathbf{x}+\frac{\mathbf{y}-\mathbf{x}}{\|\mathbf{y}-\mathbf{x}\|}t)dt\right),$ (4)
@@ 3.2 Student Model
[…]
@@
$f^{p}_{lm}(\boldsymbol{\mathbf{\mu}})=\int_{S^{2}}\text{AO}^{p}\left(\boldsymbol{\mathbf{\mu}},\boldsymbol{\mathbf{\omega}}\right)Y_{l}^{m}(\boldsymbol{\omega})d\boldsymbol{\omega},$ (14)
@@
where degdeg is the degree of SH. To model the shadowing effect caused by body articulation, we transform the point $\boldsymbol{\mathbf{\mu}}_{o}$ and the normal $\boldsymbol{\mathbf{n}}_{o}$ from the Gaussian primitive in observation space to the canonical space of each body part. Then, we query the $\hat{\text{AO}}^{p}$ for each part and multiply them together to get the final ambient occlusion:
@@ 4.4 Intrinsic Properties Comparisons
We also compare with R4D and IntrinsicAvatar for the task of intrinsic property decomposition on the RANA dataset. As shown in the Tab. 2, our method outperforms our teacher model in terms of normal consistency. This disparity becomes more visible in qualitative comparisons shown in Fig. 3: IA tends to have texture details baked into geometry, whereas our model successfully keeps meaningful wrinkles while discarding high-frequency noise during distillation, thanks to the smoothness prior from 2DGS. For albedo estimation, our student model achieves comparable accuracy to that of the teacher model. Although increasing the albedo distillation loss enables the student model to closely match the teacher’s albedo, this exact replication proves to be suboptimal for relighting under our approximated PBR pipeline for student. Instead, our objective prioritizes improved relighting performance over precise intrinsic decomposition.
@@ 5 Conclusion
Acknowledgement: Zeren Jiang is supported by Clarendon Scholarship. Shaofei Wang and Siyu Tang acknowledge SNSF grant 200021 204840.
@@
In this supplementary document, we provide additional materials to supplement our main submission. The code is available here for research purposes: github.com/jzr99/DNF-Avatar
@@ 7.5 Rendering Speed
8 Additional Discussion
Figure 8: Visualization of intra/inter-part occlusion.
8.1 Clarification for Part-wise Occlusion
Ambient occlusion is calculated on the fly during animation. The idea is that each body part is rigid, and thus we can pre-compute its occlusion probes in canonical space of each part. The pre-convolution (Eq. (9)) ensures that a single query is sufficient to obtain ambient occlusion at test time. For a single Gaussian in observation space, we transform it to the canonical space of $N_{p}$ body parts according to per-part rigid transforms. By querying partwise ambient occlusion probes in canonical space, we obtain $N_{p}$ ambient occlusion values, and the product of these values is the final ambient occlusion (Eq. (16)). We also present Fig. 8, where intra-part occlusion (calculated when Gaussian was transformed to the canonical space of its own part) captures the shadow of the wrinkle and local geometry, while inter-part occlusion (calculated when Gaussian was transformed to the canonical space of other parts) models e.g., shadow cast by the body onto the inner side of the arm. Posed body geometry can be used with Monte Carlo (MC) ray-tracing to compute shadows but it’s not real-time. Our pre-computed part-wise occlusion probes avoid MC ray-tracing, enabling real-time rendering.
8.2 Worse Albedo but Better Relighting Results
The ground-truth dataset and the teacher model both employ MC ray tracing. If we use a large loss to enforce the albedo from the teacher to the student, the final relighting results will be suboptimal since split-sum is an approximation to ray-tracing. Hence, we use a small weight for albedo distillation, which serves more like a regularization to make sure the student does not produce unreasonable albedos. Our albedos are thus optimized for split-sum and are not consistent with the ground-truth, which employs ray-tracing. On the other hand, our learned normals are less noisy than the teacher, while split-sum does not suffer from MC noise. These factors combined give us a better relighting result.
@@ 9 Limitations and Societal Impact Discussion
Moreover, our current per-scene optimization-based pipeline remains slow during training. Similar to other state-of-the-art feed-forward dynamic reconstruction methods [30, 35, 73], a promising future direction is to learn a data-driven prior for intrinsic property decomposition, enabling a feed-forward approach for animatable and relightable avatar reconstruction.